OpenStudy (anonymous):

through: (2,1) , Slope= 2

OpenStudy (ipwnbunnies):

Find an equation of a line?

OpenStudy (anonymous):

yes , Im Having Trouble On Theese Type of Problems @iPwnBunnies

OpenStudy (ipwnbunnies):

Ok, use the point-slope equation to build an equation for a line. \[y - y1 = m(x-x1)\] Where (x1,y1) is the given point and 'm' is the slope.

OpenStudy (anonymous):

so how do I set the problem up

OpenStudy (ipwnbunnies):

Plug in the numbers, then expand it, then solve for y. \[y - 1 = 2(x-2)\]

OpenStudy (anonymous):

so y=2x-3

OpenStudy (ipwnbunnies):

Yes. :)

OpenStudy (anonymous):

thats all i have to do for theese types of problems

OpenStudy (ipwnbunnies):

Ok, well, you got it.

OpenStudy (anonymous):

Thanks So Much , One More Question If I have A Problem With No Slope Then How Do I Set It Up Same Way ... Ex: Through (0,2) and (-4,1)

OpenStudy (ipwnbunnies):

You'll have to find the slope of the line formed between the two points, using slope formula. Then, choose a point, and use the point-slope equation again.

OpenStudy (ipwnbunnies):

\[slope = \frac{Y2 - Y1}{X2 -X1}\] Where (X1,Y1) is one point and (X2,Y2) is the other point.